OpenStudy (anonymous):

Algebraically determine the x-intercept and y-intercept. x/5+y/2=1

OpenStudy (watchmath):

Set \(x=0\) we have \(y/2=1\). So \(y=2\). Hence (0,2) is the y-intercept. Set \(y=0\) we have \(x/5=1\). SO \(x=5\). Hence (5,0) is the x-intercept.

OpenStudy (anonymous):

How would I write out that equation in slope-intercept form?

OpenStudy (watchmath):

You just need to isolate the \(y\). First move the \(x/5\) \(y/2=-x/5+1\) Then multiply by 2 \(y=-\frac{2}{5}x+2\)